OpenStudy (anonymous):

4(X + 6) = 100 76 19 20 95

OpenStudy (anonymous):

4(x + 6) = 100 -- distribute through the parenthesis 4x + 24 = 100 -- subtract 24 from both sides 4x = 100 - 24 4x = 76 -- now divide both sides by 4 to get x by itself 4x/4 = 76/4 x = 19

OpenStudy (anonymous):

^ That is correct.

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

19